the function f:R->R defined by f(x) = 4^(x) – 4^|x|.
Find its range?

Grade:12

1 Answers

Arun
4 years ago
Dear student
 
if x> 0
f(x) = 4^x – 4^x = 0 ( always)
 
if x
 
f(x) = 4^x – 4^-x 
 
hence now range = (- infinty, 0]